OpenStudy (anonymous):

show that e^x>1+x+(x^2/2), for all x>0 @MIT 18.02 Multiva…

OpenStudy (anonymous):

If able, write out the maclaurin series for e^x. That shows this

OpenStudy (anonymous):

depends on your definition of \[e^x\]

OpenStudy (anonymous):

what do you mean depending on the definition of e^x

OpenStudy (jamesj):

Consider the function f(x) = e^x - (x^2/2 + x + 1). Note that for x = 0, f(x) = 0. Now if you can show that the derivative of f'(x) is always positive for x > 0, that will do the trick. f'(x) = e^x - (x + 1) Still a bit of a problem. But notice now that f'(x) = 0. If we can show that (f'(x))' = f''(x) > 0 for x > 0, then that will do the trick. f''(x) = e^x - 1. As e^x > 1 for x > 0, it is indeed the case that f''(x) > 0 for x > 0. Hence f"(x) > 0 for x > 0 And therefore f(x) > 0 for x > 0. ====== It's possible now to generalize this solution to show that for \( x > 0 \) \[ e^x > \sum_{i=0}^n x^i/i! \]

OpenStudy (jamesj):

("Hence f"(x) > 0 for x > 0" should read "Hence f'(x) > 0 for x > 0")

OpenStudy (anonymous):

I really like that proof. Nice